### v1

# RoDUS: Robust Decomposition of Static and Dynamic Elements in Urban Scenes

Abstract
The task of separating dynamic objects from static environments using NeRFs has been widely studied in recent years. However, capturing large-scale scenes still poses a challenge due to their complex geometric structures and unconstrained dynamics. Without the help of 3D motion cues, previous methods often require simplified setups with slow camera motion and only a few/single dynamic actors, leading to suboptimal solutions in most urban setups. To overcome such limitations, we present RoDUS, a pipeline for decomposing static and dynamic elements in urban scenes, with thoughtfully separated NeRF models for moving and non-moving components. Our approach utilizes a robust kernel-based initialization coupled with 4D semantic information to selectively guide the learning process. This strategy enables accurate capturing of the dynamics in the scene, resulting in reduced artifacts caused by NeRF on background reconstruction, all by using self-supervision. Notably, experimental evaluations on KITTI-360 and Pandaset datasets demonstrate the effectiveness of our method in decomposing challenging urban scenes into precise static and dynamic components.
Keywords:
Neural Radiance Fields (NeRFs) Static-Dynamic Decomposition Motion Segmentation
Figure 1: RoDUS is a neural scene representation designed to decompose 4D dynamic scenes into two elements: moving foreground and static background. This decomposition remains consistent across all photometric, geometric, and semantic aspects.
1 Introduction
There have been numerous works that achieved success in various 3D representations including both implicit [23] and explicit [36, 37] scene modeling. Therefore, reconstructing and decomposing large-scale dynamic outdoor scenes is becoming a critical task for 3D scene understanding, especially for domains such as closed-loop simulation for autonomous driving [48, 54, 40]. Accurate segmentation and removal of dynamic objects are crucial as they cause floating artifacts during novel view synthesis (NVS) of background scenes due to view inconsistency. This inevitably affects other downstream applications. For instance, several works [20, 7, 12] require masking out eventual dynamic objects as they estimate camera poses. Indeed, in the case of highly dynamic scenes, without feeding the correct masks, they tend to make wrong refinements of the poses. On the other hand, dynamic objects often contain valuable information in the scene and play a main role in tasks such as surveillance and video compression [22].
However, advancements are lagging as understanding time-dependent scenes is still an ill-posed problem. This is because it demands considerable efforts to detect and treat moving components separately, while still ensuring spatio-temporal consistency compared to static scenes [30, 44]. The challenges further arise in driving settings, featuring not only large unbounded scenes, with lots of unconstrained dynamics and various illuminations but also large spatio-temporal gaps captured from the ego-vehicle. This illustrates that a majority of the background regions are captured only for a short duration, due to high occlusion caused by other dynamic objects while the ego-vehicle is also in continuous motion. Unlike previous works [28, 29, 18, 47], where scenes typically involve a single object with slow ego-camera motion, the scenarios encountered in driving sequences are less controllable.
Until now, the research community has proposed various solutions to the challenge of decoupling dynamic objects from static environments. Many of those develop NeRF-based multi-branch architectures in which only the dynamic branch (i.e. time-dependent branch) is employed for rendering dynamic objects while the background is learned using a separate static branch. Subsequently, these methods need to detect moving objects offline, so that the static model could later avoid sampling points at regions that belong to dynamic objects [27, 15, 48, 50, 53, 60, 51]. Nevertheless, they end up relying on expensive annotations for segmenting and tracking every object on the scene to enable correct model guidance. Such limitation restricts their adaptability to other datasets. Others [21, 4] learn to classify the inconsistency through a rejection factor while jointly optimizing both branches. However, NeRF’s view-dependent capacity makes it difficult to distinguish non-Lambertian but static from true dynamic objects, requiring careful and expensive parameter tuning to reach optimal separation.
In this paper, by building upon self-supervised 4D scene representation using NeRF, RoDUS (Robust Decomposition of Static and Dynamic Elements in Urban Scenes) employs a two-pathway architecture with additional sky and shadow modeling that is suitable for urban scenes. While the architecture is mainly used for representing dynamic scenes as a whole, we argue that there appears the possibility of both branches learning incorrect information (i.e. local minima), as long as the overall reconstruction satisfies the ground truth image. This deviates from the initial design intent, where the static branch is meant to learn only the static background, and vice versa. Toward this goal, we inspire from [1] by proposing a robust kernel-based strategy, thus reducing L2 sensitivity. Additionally, we introduce a view-independent semantic field to prevent the static model from expressing unwanted regions, while still being able to refine high-frequency detail. This semantic awareness is further used to guide the information flow, thus balancing the contribution of each branch. To summarize, the main contributions of our work are:
  • We demonstrate RoDUS’s capability of static-dynamic decomposition across photometric, geometric, and semantic outputs in a self-supervised manner.
  • We introduce a robust initialization scheme and semantic reasoning ability with a masking mechanism that ensures the quality of separation.
  • We conduct comprehensive comparisons on several challenging driving sequences, demonstrating our RoDUS’s performance on static-dynamic components extraction, and decoupling compared to other SOTA methods.
2 Related Work
Recently, many works have achieved promising results for high-fidelity view synthesis [2], increasing the training and inference speeds [25], and combining with semantic [9, 15, 58, 13], foundation model-guided [14, 41] reasoning to suit urban driving scenarios [33, 10]. As most of the above methods focus on static scenes, some other works have targeted non-static, transient elements from video sequences [28, 29, 18, 17, 30, 44].
NeRFs for outliers rejection. NeRF-W [21] is the pioneering work that decomposes the scene into static and transient components to reconstruct the landmark from unconstrained phototourism datasets. They propose per-frame latent embeddings and a transient head to model different lighting conditions and transient effects, thus these phenomena can be removed from the rendering function at inference. Alternatively, RobustNeRF [34] successfully rejects distractors from the scenes using a trimmed kernel to modify the photometric loss into an Iteratively Least Squares problem, they also propose patch filtering to maintain sharp reconstruction quality. Ha-NeRF [4], instead, parameterizes the kernel as learnable and uses the 2D visibility map per image as a weight function. This occlusion-aware module is proven to be more accurate at separating transient subjects.
Motion segmentation. In addition to advancements in 3D vision, there has been extensive research focusing on self-supervised motion segmentation at 2D image level. The majority of these methods rely on 2D trackers [7], optical flow patterns [3, 16, 55] and/or image warping [35] to distinguish between ego motions and object motions and segment them. These approaches have clear limitations as they focus solely on image-level processing and lack scene-level understanding. Therefore, they struggle to handle complicated static backgrounds with large camera motion [47]. Furthermore, those who follow the pipeline of first-process-then-NeRF create not only label ambiguity that affects learning but also add extra costs for separate perception processing tasks, considering that motion segmentation or flow classification themselves are already challenging.
NeRFs for dynamic scene reconstruction. To represent dynamic scenes, a direct solution is to add time as an additional input [32, 17] or learn a deformation field that maps the input coordinates into a canonical space [28, 29]. These time-dependent extensions, however, fail to accurately represent complex scenes and are unable to disentangle the dynamic components. Several works have adopted the idea in multi-branch architectures (e.g. NSFF [18], NeuralDiff [42], STaR [56], D2NeRF [47]), decomposing the scene using different radiance fields. Nonetheless, the mentioned methods lack robustness and are not tested in urban driving scenarios where it is much harder to manage the number of dynamic objects within the scene. Consequently, their performance significantly degrades due to the blurriness of dynamic objects and floating artifacts in background regions, restricting their applicability. Follow-up works exploit object-centric approaches, especially NGS [27], PNF [15], and DisCoScene [50] express dynamic scenes and target automotive data, but they are fundamentally constrained by per-scene 3D annotations including bounding boxes and instance tracking which are difficult to obtain reliably without 3D sensors and proper calibration [11].
SUDS [43], and EmerNeRF [52] are the most similar to our approach since they utilize 2D off-the-shelf models to enable object- and scene-level semantic understanding while still being able to separate dynamic objects. Our approach, however, differs from existing works: we pay more attention to the representation of each branch individually, a factor often neglected by other studies but stands advantageous in handling such challenging data. This gives us better performance in lots of downstream tasks such as background reconstruction, object removal, and motion segmentation.
3 Method
This section details RoDUS’s methodology of disentangling and composing representation in addition to its computational pipeline which is illustrated in Fig. 2. We describe in Sec. 3.1 the architecture for our scene representation. In Sec. 3.2, we introduce the semantic-radiance field used in RoDUS, how it can be rendered using differentiable volume rendering and contributes to the separation. Finally, Sec. 3.3 formulates the losses to train RoDUS as well as our proposed training strategy, which is demonstrated in Sec. 3.4.
Goals. Our main goal is to learn a scene representation of a dynamic environment (e.g. driving scene) with the ability to semantically disentangle between static components (e.g. road, building, vegetation, parked vehicles) and dynamic objects in a self-supervised manner. RoDUS takes a set of $N$ input RGB images of size $W\times H$, their associated poses, and capturing timestamps $\{I_{i}^{H\times W\times 3},p_{i},t_{i}\}_{i=1}^{N}$. Additionally, our model relies on semantic $\{S_{i}^{H\times W\times L}\}_{i=1}^{N}$, where $L$ is the number of classes, and geometric cues obtained from pre-trained 2D segmentation model and LiDAR depth measurements, respectively. At inference time, we wish to achieve rendered images for static $\{\hat{I}_{i}^{S},\hat{S}_{i}^{S}\}_{i=1}^{N}$ and dynamic $\{\hat{I}_{i}^{D},\hat{S}_{i}^{D}\}_{i=1}^{N}$ components of the scene for each timestamp independently, which can also be combined for a complete scene. As mentioned above, we study the quality of individual components rather than conforming solely to a satisfying composed result.
3.1 Scene Representation
Preliminaries. RoDUS is based on NeRF [23] which is designed to capture both geometry and view-dependent appearance of a scene from a set of RGB images. It encodes the scene within a Multi-Layer Perceptron (MLP) and outputs color $\mathbf{c}(\mathbf{x},\mathbf{d})\in\mathbb{R}^{3}$ and density $\sigma(\mathbf{x})\in\mathbb{R}$ at any query position $\mathbf{x}\in\mathbb{R}^{3}$ along with its viewing direction $\mathbf{d}\in\mathbb{R}^{3}$. During training, using intrinsic parameters and poses, every pixel of the image undergoes a ray marching procedure $\mathbf{r}$ to sample inputs for the MLP. The model parameters are optimized to predict the colors rendered along the ray using the mean squared error: $\mathcal{L}=\|\hat{\mathbf{C}}_{i}(\mathbf{r})-\mathbf{C}_{i}(\mathbf{r})\|_{2}^{2}$.
Architecture. Inspired by [21, 18, 47], we factorize the scene into two elements, each learned through a separate pathway: one for moving components and the other for non-moving background. Each branch is modeled independently using its own multi-level hash grid representation and MLPs. This approach enables the concurrent learning of both static and dynamic aspects of the scene.
The static branch whose mission is to learn all non-moving parts in the scene, uses a 3D hash grid $\mathcal{H}^{S}$, analogous to the one used in [25]. To model lighting variations, we also condition on a per-frame latent embedding $l_{a}$ [21]:
$(\sigma^{S},\mathbf{c}^{S},\mathbf{s}^{S})=\mathcal{F}^{S}_{\Theta}(\mathcal{H}^{S}(\mathbf{x}),\gamma(\mathbf{d}),l_{a}).$ (1)
As the dynamic branch allows its outputs to depend on time, we employ a 4D hash grid $\mathcal{H}^{D}$ with one additional normalized timestamp $t$ as input. The dynamic field additionally incorporates a shadow head [47] designed to learn a shadow ratio $\rho\in[0,1]$. This scalar is later used to scale down the static radiance values $\mathbf{c}^{S}$ in Eq. 3:
$\begin{split}(\sigma^{D},\mathbf{c}^{D},\mathbf{s}^{D})&=\mathcal{F}^{D}_{\Theta}(\mathcal{H}^{D}(\mathbf{x},t),\gamma(\mathbf{d})),\\ \rho&=\mathcal{F}^{\rho}_{\Theta}(\mathcal{H}^{D}(\mathbf{x},t)).\end{split}$ (2)
Figure 2: RoDUS’s architecture. Our model comprises two separate branches that take as input sampled positions $\mathbf{x}$, their viewing direction $\mathbf{d}$, and their timestamp $t$ for the dynamic branch, and generate outputs for every query coordinate. Each branch is represented by a separate hash grid and the MLP-based neural function, which predicts colors, densities, and semantics. The rendered static image is used to calculate IRLS map $\omega(\epsilon)$ during the robust initialization step (Sec. 3.3), while dynamic semantic outputs are passed through a “foreground-only mask” to prevent over-explaining background regions (Sec. 3.2).
Rendering. We calculate the color $\hat{C}$ for each camera ray $\mathbf{r}$ by integrating all the sampled points along the ray:
$\hat{C}(\mathbf{r})=\sum_{i=1}^{K}T_{i}\left(\alpha_{i}^{S}(1-\rho_{i})\mathbf{c}_{i}^{S}+\alpha_{i}^{D}\mathbf{c}_{i}^{D}\right),$ (3)
where $T_{i}=\exp\left(-\sum_{j=1}^{i-1}(\sigma_{j}^{S}+\sigma_{j}^{D})\delta_{j}\right)$ is the accumulated transmittance and $\alpha_{i}^{\xi}=1-\exp(-\sigma_{i}^{\xi}\delta_{i})$ with $\xi\in\{S,D\}$ are alpha values. Given the outputs of each branch, we can render static and dynamic maps separately.
3.2 Enabling Semantic Awareness
NeRF provides low-level implicit representations of geometry and radiance but lacks a higher-level understanding of the scene. It is shown in [34] that relying only on view-dependent radiance and sensitive mean square error makes the model prone to outliers and over-fit observations. On the other hand, semantic labels can be conceptualized as a view-invariant function [59], mapping world coordinates $\mathbf{x}$ to a distribution over semantic labels $\mathbf{s}(\mathbf{x})\in\mathbb{R}^{L}$. As a result, we choose to integrate a semantic head to predict the semantic class for each point, similar to the color head. Subsequently, the same rendering formula is applied for semantics to achieve 2D semantic segmentation maps:
$\hat{S}(\mathbf{r})=\sum_{i=1}^{K}T_{i}\left(\alpha_{i}^{S}\mathbf{s}_{i}^{S}+\alpha_{i}^{D}\mathbf{s}_{i}^{D}\right).$ (4)
The use of semantic supervision in RoDUS serves two main purposes. First, it provides assumptions for multiple classes, forming the basis for the loss formulations (refer to Sec. 3.3) that contribute to enhancing rendering quality. Second, it enables the model to capture detailed scene layout and object boundary information with a high level of precision thereby assisting in decomposition.
Figure 3: (a) We enforce a “foreground-only mask” to the dynamic semantic head. In the forward pass (black arrows), the mask prevents the dynamic branch from outputting pixels that do not belong to foreground classes. Similarly, in the backward pass (orange arrows), it restricts the dynamic field from learning background pixels, which may primarily come from noisy annotations. (b) Predictions generated by 2D segmentation model are noisy and inconsistent between views. (c) Since the dynamic field includes a temporal dimension, it ends up learning these conflicting labels, satisfying the overall loss. (d) Therefore, our proposed mask is used to tackle the problem.
One point to consider is that supervision for segmentation with noisy pseudo labels should not impact the underlying geometry. Otherwise, the model can still satisfy conflicting labels by changing the dynamic weights, resulting in the wrong geometry. Previous works [9, 39] stop gradients from the semantic head back to the densities to solve this problem. However, we believe that semantic information plays an important role in correcting the separation. We instead limit the dynamic model from contributing static class prediction to the final results by putting a mask in the last layer of its semantic head. This masking strategy can be viewed as a form of gradient stopping the learning but in a selective way as illustrated in Fig. 3. Note that we do not have motion information for each instance, we only leverage the fact that certain classes are potentially movable to penalize the dynamic model from predicting background (non-movable) samples.
We apply softmax normalization at each query point individually before ray integration, as introduced in [39]. This is done due to the unbounded nature of logits, ensuring that the semantic head can not generate excessively high scores in regions with low density to overcome high-density areas. Remarkably, this not only improves the stability of rendering semantics but also prevents half-precision floating point (float16) overflowing when working with frameworks such as tiny-cuda-nn [24].
(a) Ground-truth
(b) $\mathcal{L}_{robust}$
(c) $\mathcal{L}_{robust}+\mathcal{L}_{sky}$
(d) $\mathcal{L}_{robust}+\mathcal{L}_{road}$
(e) $\mathcal{L}_{robust}+\mathcal{L}_{sky}+\mathcal{L}_{road}$
Figure 4: (b) The use of a trimmed kernel helps completely remove moving cars in the scene. (c, d, e) As the sky and road regions are badly reconstructed, applying class constraints can aid the learning process of the static field.
3.3 Loss Formulation
Reconstruction losses. Consistent with other NeRF works, we incorporate both L2 and DSSIM [46] losses to supervise the pixel color $\mathcal{L}_{rgb}$. In cases where LiDAR data is available, depth supervision [8, 33] loss $\mathcal{L}_{d}=\|\hat{d}_{i}(\mathbf{r})-d_{i}(\mathbf{r})\|_{2}^{2}$ is considered to enhance the reconstruction accuracy. In addition, cross-entropy [59] loss is applied to provide supervision on the generated semantic labels derived from pseudo-ground truth segmentation, expressed as $\mathcal{L}_{sem}=CE(\hat{S}(\mathbf{r}),S(\mathbf{r}))$.
Robust loss. Our objective is to reduce the sensitivity of the static model to dynamic objects by incorporating supervision through a robust photometric loss. This loss can be formulated as an Iteratively Reweighted Least Squares (IRLS) problem:
$\mathcal{L}_{robust}=\kappa\left(\|\hat{C}^{S}(\mathbf{r})-C(\mathbf{r})\|_{2}\right)=\omega(\epsilon(\mathbf{r}))\cdot\|\hat{C}^{S}(\mathbf{r})-C(\mathbf{r})\|_{2}^{2},$ (5)
where $\epsilon(\mathbf{r})=\|\hat{C}^{S}(\mathbf{r})-C(\mathbf{r})\|_{2}$ is the error residuals and $\omega(\epsilon)=\epsilon^{-1}\cdot\partial\kappa(\epsilon)/\partial\epsilon$ is the weight function of the residuals at the previous iteration. Among the family of kernels [1], we follow Sabour et al. [34] and choose a trimmed estimator [6] which classifies outliers based on a percentile threshold $\mathcal{T}_{\epsilon}$ calculated from the whole batch of rays. This choice ensures outliers are completely removed (i.e. hard weights). Furthermore, as every object has neighbor connections, they apply a spatial filter $\mathcal{F}$ on the residuals map to capture the smoothness, ensuring local support of the kernel:
$\omega(\epsilon(\mathbf{r}))=\mathcal{F}\left[\epsilon(\mathbf{r})\leq\mathcal{T}_{\epsilon}\right],$ (6)
where $\mathcal{F}$ consists of a box filter to diffuse inlier/outlier labels and a sub-patch level consistency filter that takes into account the rejection behavior of its surrounding neighborhoods. This filter aims to remove high-frequency details from being classified as outliers and allows them to be captured by the NeRF models.
Sky loss. Outdoor scenes contain sky regions where rays never intersect any opaque surfaces. We detect pixels that belong to the sky and force the densities to be zeros along the ray, excluding the last sample:
$\mathcal{L}_{sky}=\mathbb{E}\left[\frac{1}{K-1}\sum_{i=1}^{K-1}\sigma_{i}+\frac{1}{\sigma_{K}}\right]_{\mathbf{r}\in\mathcal{R}_{sky}}.$ (7)
Planar regularization loss. We observe that the road surface is challenging to model as it has low textures and is often occluded by moving vehicles. These areas mostly suffer from poor geometry prediction (Fig. 4). To address this issue, we rely on the work of Wang et al. [45], perform Singular Values Decomposition (SVD) on the set of centered points projected from patches that are assigned with the road label, and minimize the smallest singular value $\sigma_{3}$:
$\mathcal{L}_{road}=\sigma_{3}(\mathcal{R}_{road}^{patch}).$ (8)
Dynamic regularization loss. We use the factorization $\mathcal{L}_{\mathcal{H}}$ and sparsity $\mathcal{L}_{\sigma^{D}},\mathcal{L}_{\rho}$ losses that have been proposed in [56, 47, 38] to enforce the static branch to explain most of the static structure as possible and only express the moving objects and/or shadows using the dynamic branch.
Our final loss results on the weighted sum of all our individual losses:
$\begin{split}\mathcal{L}_{total}&=\underbrace{\mathcal{L}_{rgb}+\lambda_{d}\mathcal{L}_{d}+\lambda_{sem}\mathcal{L}_{sem}}_{reconstruction}\\ &+\underbrace{\mathcal{L}_{robust}+\lambda_{sky}\mathcal{L}_{sky}+\lambda_{road}\mathcal{L}_{road}}_{static\ regularization}+\underbrace{\lambda_{\sigma^{D}}\mathcal{L}_{\sigma^{D}}+\lambda_{\rho}\mathcal{L}_{\rho}+\lambda_{\mathcal{H}}\mathcal{L}_{\mathcal{H}}}_{dynamic\ regularization}\end{split}$ (9)
3.4 Training strategy
The use of robust estimation plays an important role in rejecting inconsistency (dynamic objects in our case) and guiding the static branch without any prior knowledge. However, we observe that using a binary kernel alone with a fixed threshold, after reaching a certain point throughout the training, the robust loss does not contribute to the overall loss anymore. We refer to this phenomenon as “kernel saturation”, this causes several parts of the scene not to be learned completely. Adjusting the threshold $\mathcal{T}_{\epsilon}$ seems to be a trivial solution, but given the fact that outdoor scenes are more complicated to handle, there is always a trade-off between the number of distractors removed and the static scene quality. Particularly, high-frequency but static details, equivalently appear as large residuals are considered as outliers, while low-frequency but moving regions such as shadows (dark) which mostly share a similar color with the road (gray) are uncontrollably included as inliers. Figure 4 portrays this, as there are still parts of the scene (e.g. the truck on the right) that are not fully learned by the model.
In contrast, relying solely on reconstruction losses leads to several local optima during optimization due to the entanglement of geometry and appearance. We propose using the robust loss as an initialization step for the static model to learn the basic overall structures of the scene before delving into complete scene understanding. The initialization step guides the network to prioritize consistent (or confident) regions that further help stabilize the training process. Therefore, we start the model training with the robust loss (Eq. 5) and sky loss (Eq. 7) enabled for the first epoch, and subsequently enable the remaining losses. This serves as a foundation for disentangling the scene in later training stages.
GT
$\hat{I}$
$\hat{d}$
$\hat{I}^{S}$
$\hat{d}^{S}$
$\hat{I}^{D}$
$\hat{S}$
$\hat{S}^{S}$
Figure 5: We demonstrate RoDUS’s capabilities on multiple tasks. Notice that we successfully retain separation consistencies between all photometric, geometric, and semantic outputs. Legend: GT: Ground truth, $\hat{I}$: Composed RGB, $\hat{d}$: Composed depth, $I^{S}$: Static RGB, $\hat{d}^{S}$: Static depth, $\hat{I}^{D}$: Dynamic RGB (blended with a green background), $\hat{S}$: Composed segmentation map, $\hat{S}^{S}$: Static segmentation map, $M^{D}$: Motion segmentation.
$M^{D}$
4 Evaluation
4.1 Experimental Setup
Implementation details. We build RoDUS using PyTorch [31]. Our implementation is based on tiny-cuda-nn [24] from iNGP [25], and proposal sampling strategy inspired from MipNeRF360 [2]. To generate pseudo-ground truth for semantic segmentation, we use the Mask2Former [5] model pre-trained on the Mapillary Vistas [26] dataset.
Datasets. We evaluate our system on two challenging urban driving datasets, KITTI-360 [19] and Pandaset [49], both featuring dynamic objects in driving scenarios. In Pandaset, we utilize three front cameras and one back camera, while in KITTI-360, we use all two perspective cameras along with side fish-eye cameras. For evaluation, we rely on provided ground-truth instance masks (KITTI-360)/projected 3D bounding boxes (Pandaset) with motion labels to identify true dynamic objects. We refer to the supplementary for additional details on the selection of the training sequences and data pre-processing.
[…]
Table 1: Quantitative comparisons with SoTA methods on KITTI-360 and Pandaset.
Image Reconstruction Static NVS Motion segmentation
[…]
KITTI-360 RobustNeRF [${}^{{\dagger}}$ 19.50 0.722 0.282 - - -
[…]
SUDS [43] 28.16 0.836 0.175 18.81 0.687 0.298 93.89 5.35 9.34
EmerNeRF [52] 25.49 0.798 0.181 20.59 0.733 0.249 57.74 12.56 19.92
RoDUS (ours) 26.47 0.857 0.148 23.43 0.802 0.181 79.45 64.84 76.52
Pandaset RobustNeRF [${}^{{\dagger}}$ 23.41 0.764 0.197 - - -
[…]
SUDS [43] 30.28 0.851 0.138 25.34 0.788 0.182 - - -
EmerNeRF [52] 29.51 0.841 0.155 27.47 0.811 0.166 - - -
RoDUS (ours) 28.39 0.877 0.135 25.83 0.825 0.156 - - -
${}^{{\dagger}}$ As RobustNeRF rejects distractors during training, full image reconstruction metrics are not included.
GT
RobustNeRF [34]
N/AN/A
[…]
SUDS [43]
EmerNeRF [52]
RoDUS (ours)
Composed RGB
Static RGB
Dynamic RGB
[…]
4.2 Static-Dynamic Decomposition
[…]
4.3 Dynamic Objects Segmentation
[…]
GT image Refer to caption Refer to caption Refer to caption
GT mask Refer to caption Refer to caption Refer to caption
[…]
SUDS [43] Refer to caption Refer to caption Refer to caption
EmerNeRF [52] Refer to caption Refer to caption Refer to caption
RoDUS (ours) Refer to caption Refer to caption Refer to caption
(a) (b) (c)
Figure 7: Motion segmentation qualitative results comparison on KITTI-360 before thresholding. We present results in three scenarios: (a) The ego-car stops in front of an intersection; (b) Vehicles move in the same direction as the ego-car; and (c) Vehicles move in the opposite direction toward the ego-car.
4.4 Ablation Study
We analyze our design choices in Sec. 4.4. The use of road, sky regularization, and robust initialization terms support the geometry, especially in regions occluded by dynamic objects, thus improving the overall quality. Please refer to Fig. 4 and Fig. 8 for visual comparisons.
[…]
Refer to caption w/o $\mathcal{L}_{road}$ Refer to caption w/o $\mathcal{L}_{robust}$ Refer to caption Full model static RGB Refer to caption Full model static depth Figure 8: Static depth map ablation
4.5 Limitation
Our approach can’t handle areas that lack observations or have been highly occluded most of the time within the sequence. This is particularly notable by large or slow-moving objects, illustrated in Fig. 9. Exploring a model that is capable of detecting and inpainting these missing regions based on prior knowledge will be interesting as a future research direction.
GT
$\hat{I}^{S}$
$\hat{S}^{S}$
$\hat{I}$
$\hat{I}^{D}$
$M^{D}$
Figure 9: Failure case. The static field is unable to recover regions that are highly occluded by the truck due to its lack of generalizing ability to unseen backgrounds. On the other hand, the dynamic field doesn’t support object-level connection, leading to only half of the truck being visible in dynamic outputs.
5 Conclusion
In this paper, we presented RoDUS, a method for disentangling dynamic scenes using neural representation. Our approach effectively separates moving components from the scene without the need for manual labeling. To enhance disentangling efficiency, we proposed a robust training strategy and incorporated semantic reasoning, resulting in high-quality background view synthesis and accurate segmentation. We conducted extensive experiments on challenging autonomous driving datasets, showcasing that RoDUS surpasses current state-of-the-art NeRF frameworks.
Appendix 0.A Implementation Details
0.A.1 Data Processing
In our experiments, we downsample Pandaset images by a factor of 4 ($480\times 270$) and KITTI-360 by a factor of 2 ($704\times 188$). The selected scenes from KITTI-360 are summarized in Tab. 3. For Pandaset, we used sequences 023, 042, and 043. We specifically selected those scenes because they feature multiple dynamic objects while providing sufficient views of the background regions.
Table 3: Selected KITTI-360 sequences for training, with skip step of 2.
Sequence Scene Start Frame End Frame #Frames per camera
0 0002 4970 5230 130
1 0003 100 330 115
2 0007 2680 2930 125
3 0010 910 1160 125
For semantic awareness, the classes we assume to be dynamic and keep the contribution in the dynamic semantic head are: “Person”, “Rider”, “Car”, “Truck”, “Bus”, “Train”, “Motorcycle”, “Bicycle”, “Bicyclist”, “Motorcyclist”, and “Unknown vehicle”.
0.A.2 Architecture
We adopt a proposal-based coarse-to-fine sampling technique[2] using two proposal models with the number of samples being $[128,96]$ for the proposal models, and 48 for the final model. Our hash grids have the size of $2^{19}$ for $16$ levels with the base resolution and finest resolution of 16 and 2048 respectively. Additionally, our dynamic hash grid allows 4D inputs, which means that the intermediate feature vector is computed through a quadrilinear interpolation of 16 grid point vectors instead of the 8 used for the 3D hash grid. The static branch of our model is equipped with a 16-dimension per-frame embedding to adapt to varying lighting conditions.
0.A.3 Losses
The individual components used in the dynamic regularization loss are:
Static-dynamic entropy factorization loss. We assume that any point in space can either belong to a static or dynamic object at one time, but not both, and encourage the transparency ratio at each point to be either 0 or 1, aiding the model in achieving a less entangled decomposition:
$\mathcal{L}_{\mathcal{H}}=\mathbb{E}\left[\mathcal{H}\left(\frac{\alpha^{D}}{\alpha^{S}+\alpha^{D}}\right)(\alpha^{S}+\alpha^{D})\right],$ (10)
where $\mathcal{H}(x)=-(x\cdot\log(x)+(1-x)\cdot\log(1-x))$ is the binary entropy function and the term is weighted by the total transparency $(\alpha^{S}+\alpha^{D})$ so that a point is allowed to have an empty density, as done in [56].
Dynamic sparsity losses. The sparsity terms take as input densities $\sigma^{D}$ and shadow $\rho$ decoded from the dynamic field respectively and encourage those values to be sparse [47, 38]:
$\mathcal{L}_{\sigma^{D}}=\mathbb{E}\left[\frac{1}{K}\sum_{i=1}^{K}\sigma_{i}^{D}\right]_{\mathbf{r}\in\mathcal{R}}~{}~{}~{}\text{and}~{}~{}~{}\mathcal{L}_{\rho}=\mathbb{E}\left[\frac{1}{K}\sum_{i=1}^{K}T_{i}\alpha_{i}\rho_{i}^{2}\right]_{\mathbf{r}\in\mathcal{R}}.$ (11)
0.A.4 Training
We trained RoDUS for 50 epochs using the Adam optimizer with a batch size of 200. The learning rate was set to $0.01$, and we used cosine decay to decrease it to $0.0001$. The rays are grouped in $15\times 15$ patches to suit DSSIM, planar regularization, and robust losses. Section 0.A.4 summarizes the hyperparameters used for the training of RoDUS, we do not require manual tuning for each scene, but we believe that such tuning could potentially further enhance performance.
Table 4: Training hyperparameters
Parameter Value
$\lambda_{d}$ 0.5
$\lambda_{sem}$ 0.1
$\lambda_{sky}$ 0.03
$\lambda_{road}$ 0.1
$\lambda_{\sigma^{D}}$ 0.05
$\lambda_{\rho}$ 0.3
$\lambda_{\mathcal{H}}$ 0.01
[…]
Appendix 0.B Additional Results
Fig. 10 and Fig. 11 show additional qualitative comparisons between our RoDUS and previous methods in static-dynamic decomposition and scene reconstruction. Fig. 12 displays RoDUS’s performance on other different scenes, demonstrating that RoDUS’s dynamic branch is able to classify moving objects from parked vehicles or other static instances. Furthermore, RoDUS can be trained from noisy semantic annotations, leading to a scene-consistent semantic representation. A combination of multiple modalities also supports the model in reconstructing and segmenting thin structures and objects (e.g. light poles, and traffic signs).
GT
RobustNeRF [34]
N/AN/A
[…]
SUDS [43]
EmerNeRF [52]
Figure 10: Additional qualitative comparisons on Pandaset.
RoDUS (ours)
Composed RGB
Static RGB
Dynamic RGB
GT Refer to caption
RobustNeRF [34]
N/A
N/A
[…]
SUDS [43] Refer to caption Refer to caption Refer to caption
EmerNeRF [52] Refer to caption Refer to caption Refer to caption
RoDUS (ours) Refer to caption Refer to caption Refer to caption
GT image Refer to caption
RobustNeRF [34]
N/A
N/A
[…]
SUDS [43] Refer to caption Refer to caption Refer to caption
EmerNeRF [52] Refer to caption Refer to caption Refer to caption
RoDUS (ours) Refer to caption Refer to caption Refer to caption
Compose RGB Static RGB Dynamic RGB
Figure 11: Additional qualitative comparisons on KITTI-360.
GT
$\hat{I}$
$\hat{d}$
$\hat{I}^{S}$
$\hat{d}^{S}$
$\hat{I}^{D}$
$\hat{S}$
$\hat{S}^{S}$
Figure 12: Additional results on RoDUS’s capability.
$M^{D}$
Appendix 0.C Additional Ablation and Analysis
0.C.1 Robust Kernel Threshold
We explain the choice of percentile threshold $\mathcal{T}_{\epsilon}$ used in RobustNeRF [34] and report qualitative results in Fig. 13. With $\mathcal{T}_{\epsilon}=0.7$, there are still a lot of structures that could not be learned (e.g. the far away building), when we increase $\mathcal{T}_{\epsilon}$ to $0.9$, the shadow regions started to appear. As a result, we choose $\mathcal{T}_{\epsilon}=0.75$ as it yields the best visual quality with the least artifacts. Correspondingly, the kernel saturates as we observed no improvement as the training continues. Notice that the model could not learn the road marking in any case.
GT
RoDUS’s static RGB
RobustNeRF ($\mathcal{T}_{\epsilon}=0.7$)
RobustNeRF ($\mathcal{T}_{\epsilon}=0.9$)
RobustNeRF ($\mathcal{T}_{\epsilon}=0.75$)
Figure 13: Qualitative comparison on static RGB image with different threshold values.
0.C.2 Shadow Handling
Shadows cast by dynamic objects are also time-varying and modeled separately in our model. Fig. 14 shows that RoDUS successfully decouples the shadows from the scene.
GT Refer to caption Refer to caption Refer to caption
Shadow mask Refer to caption Refer to caption Refer to caption
Figure 14: We obtain the shadow mask by applying volumetric rendering on the shadow ratio $\rho$, demonstrating that the model is learning the correct shadow regions.
Appendix 0.D Discussions
0.D.1 Ethical Concern
Our method mainly serves the purpose of removing dynamic vehicles from the scenes. While such methods hold promise in augmenting training environments for autonomous vehicles by simulating diverse scenarios, considering the act of modifying real-world driving data for inappropriate motives (e.g. erasing actors, generating fake content) is not encouraged. Altering records for any purpose carries the risk of distorting the true nature of driving events and the credibility of data-driven insights. In contrast, we would be delighted to see directions that focus on detecting image forgery for such methods. We believe this area of study could hold significant potential for advancements in the field, providing opportunities to enhance the robustness and security of digital content.
0.D.2 Limitation
Apart from the main limitation mentioned in the paper, we also encountered several problems that are related to most NeRF-based methods in general: RoDUS demands precise camera poses, which require multiple sources of odometry when working with outdoor, high-dynamic scenes. Furthermore, for RoDUS to yield good reconstruction, it relies on a substantial number of input images from multiple views, making it less effective in few-shot settings.

### v2

@@
 Our approach utilizes a robust kernel-based initialization coupled with 4D semantic information to selectively guide the learning process.
-This strategy enables accurate capturing of the dynamics in the scene, resulting in reduced artifacts caused by NeRF on background reconstruction, all by using self-supervision.
+This strategy enables accurate capturing of the dynamics in the scene, resulting in reduced floating artifacts in the reconstructed background, all by using self-supervision.
 Notably, experimental evaluations on KITTI-360 and Pandaset datasets demonstrate the effectiveness of our method in decomposing challenging urban scenes into precise static and dynamic components.
@@
 Therefore, reconstructing and decomposing large-scale dynamic outdoor scenes is becoming a critical task for 3D scene understanding, especially for domains such as closed-loop simulation for autonomous driving [48, 54, 40].
-Accurate segmentation and removal of dynamic objects are crucial as they cause floating artifacts during novel view synthesis (NVS) of background scenes due to view inconsistency.
+Accurate segmentation and removal of dynamic objects are crucial as they are the main reason for artifacts during novel view synthesis (NVS) of background regions due to view inconsistency.
 This inevitably affects other downstream applications.
 For instance, several works [20, 7, 12] require masking out eventual dynamic objects as they estimate camera poses.
-Indeed, in the case of highly dynamic scenes, without feeding the correct masks, they tend to make wrong refinements of the poses.
+Indeed, in highly dynamic scenes, without feeding the correct masks, they tend to make wrong refinements of the poses.
 On the other hand, dynamic objects often contain valuable information in the scene and play a main role in tasks such as surveillance and video compression [22].
-However, advancements are lagging as understanding time-dependent scenes is still an ill-posed problem.
-This is because it demands considerable efforts to detect and treat moving components separately, while still ensuring spatio-temporal consistency compared to static scenes [30, 44].
-The challenges further arise in driving settings, featuring not only large unbounded scenes, with lots of unconstrained dynamics and various illuminations but also large spatio-temporal gaps captured from the ego-vehicle.
-This illustrates that a majority of the background regions are captured only for a short duration, due to high occlusion caused by other dynamic objects while the ego-vehicle is also in continuous motion.
-Unlike previous works [28, 29, 18, 47], where scenes typically involve a single object with slow ego-camera motion, the scenarios encountered in driving sequences are less controllable.
+However, advancements are lagging as understanding time-dependent scenes is an ill-posed problem.
+This is because it demands considerable efforts to detect and treat moving components separately while ensuring spatiotemporal consistency compared to static scenes [30, 44].
+The challenges further arise in driving settings, featuring not only large unbounded scenes, with lots of unconstrained dynamics and various illuminations but also large temporal gaps captured from the ego-vehicle.
+This demonstrates that a majority of the background regions are captured only for a short duration, due to high occlusions caused by other dynamic objects while the ego-vehicle is also in continuous motion.
+Unlike previous works [28, 29, 18, 47], where scenes typically involve a single object with slow ego-motion, the scenarios encountered in driving sequences are less controllable.
 Until now, the research community has proposed various solutions to the challenge of decoupling dynamic objects from static environments.
-Many of those develop NeRF-based multi-branch architectures in which only the dynamic branch (i.e.
-time-dependent branch) is employed for rendering dynamic objects while the background is learned using a separate static branch.
+Many develop NeRF-based multi-branch architectures in which only the dynamic (i.e.
+time-dependent) model is employed for rendering dynamic objects while the background is learned using a separate static model.
 Subsequently, these methods need to detect moving objects offline, so that the static model could later avoid sampling points at regions that belong to dynamic objects [27, 15, 48, 50, 53, 60, 51].
-Nevertheless, they end up relying on expensive annotations for segmenting and tracking every object on the scene to enable correct model guidance.
-Such limitation restricts their adaptability to other datasets.
+Nevertheless, they end up relying on expensive annotations for segmenting and tracking every object in the scene to enable correct guidance.
+Such limitation restricts their adaptability to many datasets.
 Others [21, 4] learn to classify the inconsistency through a rejection factor while jointly optimizing both branches.
 However, NeRF’s view-dependent capacity makes it difficult to distinguish non-Lambertian but static from true dynamic objects, requiring careful and expensive parameter tuning to reach optimal separation.
-In this paper, by building upon self-supervised 4D scene representation using NeRF, RoDUS (Robust Decomposition of Static and Dynamic Elements in Urban Scenes) employs a two-pathway architecture with additional sky and shadow modeling that is suitable for urban scenes.
-While the architecture is mainly used for representing dynamic scenes as a whole, we argue that there appears the possibility of both branches learning incorrect information (i.e.
+In this paper, by building upon self-supervised 4D scene representation using NeRF, RoDUS (Robust Decomposition of Static and Dynamic Elements in Urban Scenes) employs a two-pathway architecture with additional sky, road, and shadow modeling that is suitable for urban scenes.
+While the architecture has been mainly used for representing dynamic scenes as a whole, we argue that it appears a possibility of both branches learning incorrect information (i.e.
 local minima), as long as the overall reconstruction satisfies the ground truth image.
-This deviates from the initial design intent, where the static branch is meant to learn only the static background, and vice versa.
-Toward this goal, we inspire from [1] by proposing a robust kernel-based strategy, thus reducing L2 sensitivity.
-Additionally, we introduce a view-independent semantic field to prevent the static model from expressing unwanted regions, while still being able to refine high-frequency detail.
+This deviates from the initial design intent, where the static model is meant to learn only the static background, and vice versa.
+Toward this goal, we draw inspiration from [1] by proposing a robust kernel-based strategy, thus reducing L2 sensitivity on the static model.
+Additionally, we introduce a view-independent semantic field to prevent the dynamic model from expressing unwanted regions, while still being able to refine high-frequency detail.
 This semantic awareness is further used to guide the information flow, thus balancing the contribution of each branch.
@@
   • We introduce a robust initialization scheme and semantic reasoning ability with a masking mechanism that ensures the quality of separation.
-  • We conduct comprehensive comparisons on several challenging driving sequences, demonstrating our RoDUS’s performance on static-dynamic components extraction, and decoupling compared to other SOTA methods.
+  • We conduct comprehensive comparisons on several challenging driving sequences, indicating our RoDUS’s performance on static-dynamic components extraction, and decoupling compared to other SoTA methods.
 2 Related Work
-Recently, many works have achieved promising results for high-fidelity view synthesis [2], increasing the training and inference speeds [25], and combining with semantic [9, 15, 58, 13], foundation model-guided [14, 41] reasoning to suit urban driving scenarios [33, 10].
+Recently, many works have achieved promising results in high-fidelity view synthesis [2] and combining with semantic [9, 15, 58, 13], foundation model-guided [14, 41] reasoning to better accommodate urban driving scenarios [33, 10].
 As most of the above methods focus on static scenes, some other works have targeted non-static, transient elements from video sequences [28, 29, 18, 17, 30, 44].
 NeRFs for outliers rejection.
-NeRF-W [21] is the pioneering work that decomposes the scene into static and transient components to reconstruct the landmark from unconstrained phototourism datasets.
+NeRF-W [21] is the pioneering work that decomposes the scene into static and transient components to reconstruct landmarks from unconstrained phototourism datasets.
 They propose per-frame latent embeddings and a transient head to model different lighting conditions and transient effects, thus these phenomena can be removed from the rendering function at inference.
-Alternatively, RobustNeRF [34] successfully rejects distractors from the scenes using a trimmed kernel to modify the photometric loss into an Iteratively Least Squares problem, they also propose patch filtering to maintain sharp reconstruction quality.
-Ha-NeRF [4], instead, parameterizes the kernel as learnable and uses the 2D visibility map per image as a weight function.
-This occlusion-aware module is proven to be more accurate at separating transient subjects.
+Alternatively, RobustNeRF [34] successfully rejects distractors from the scenes using a trimmed kernel to modify the photometric loss into an Iteratively Least Squares problem, they also include patch filtering to maintain sharp reconstruction quality.
+Ha-NeRF [4], instead, parameterizes the kernel as learnable and uses 2D visibility map per image as a weight function.
+This occlusion-aware module is proven to be more accurate at separating transients.
 Motion segmentation.
@@
 These approaches have clear limitations as they focus solely on image-level processing and lack scene-level understanding.
-Therefore, they struggle to handle complicated static backgrounds with large camera motion [47].
+Therefore, they struggle to handle complicated scenes with large camera motion [47].
 Furthermore, those who follow the pipeline of first-process-then-NeRF create not only label ambiguity that affects learning but also add extra costs for separate perception processing tasks, considering that motion segmentation or flow classification themselves are already challenging.
@@
 Follow-up works exploit object-centric approaches, especially NGS [27], PNF [15], and DisCoScene [50] express dynamic scenes and target automotive data, but they are fundamentally constrained by per-scene 3D annotations including bounding boxes and instance tracking which are difficult to obtain reliably without 3D sensors and proper calibration [11].
-SUDS [43], and EmerNeRF [52] are the most similar to our approach since they utilize 2D off-the-shelf models to enable object- and scene-level semantic understanding while still being able to separate dynamic objects.
-Our approach, however, differs from existing works: we pay more attention to the representation of each branch individually, a factor often neglected by other studies but stands advantageous in handling such challenging data.
-This gives us better performance in lots of downstream tasks such as background reconstruction, object removal, and motion segmentation.
+SUDS [43] and EmerNeRF [52] are the most similar to our approach since they utilize 2D off-the-shelf models to enable object-/scene-level semantic understanding while still being able to separate dynamic objects.
+However, our approach differs from existing works in that we place greater emphasis on the representation of each branch individually, a factor often overlooked by other studies.
+This focus proves advantageous in understanding complex scenes.
+Specifically, our dynamic model is designed to learn only the true dynamic instances rather than capturing variations in illumination or pixel changes caused by ego-motion.
 3 Method
-This section details RoDUS’s methodology of disentangling and composing representation in addition to its computational pipeline which is illustrated in Fig.
+This section details RoDUS’s methodology of disentangling and composing representation with its computational pipeline illustrated in Fig.
 2.
-We describe in Sec.
-3.1 the architecture for our scene representation.
-In Sec.
-3.2, we introduce the semantic-radiance field used in RoDUS, how it can be rendered using differentiable volume rendering and contributes to the separation.
-Finally, Sec.
-3.3 formulates the losses to train RoDUS as well as our proposed training strategy, which is demonstrated in Sec.
+Sec.
+3.1 describes the architecture for our scene representation.
+Sec.
+3.2 introduces the semantic-radiance field used in RoDUS, how it can be rendered using differentiable volume rendering and contributes to the separation.
+Finally, we formulate the losses to train RoDUS in Sec.
+3.3 as well as our proposed training strategy in Sec.
 3.4.
@@
 Our main goal is to learn a scene representation of a dynamic environment (e.g.
-driving scene) with the ability to semantically disentangle between static components (e.g.
+a driving scene) with the ability to semantically disentangle between static background (e.g.
 road, building, vegetation, parked vehicles) and dynamic objects in a self-supervised manner.
-RoDUS takes a set of $N$ input RGB images of size $W\times H$, their associated poses, and capturing timestamps $\{I_{i}^{H\times W\times 3},p_{i},t_{i}\}_{i=1}^{N}$.
-Additionally, our model relies on semantic $\{S_{i}^{H\times W\times L}\}_{i=1}^{N}$, where $L$ is the number of classes, and geometric cues obtained from pre-trained 2D segmentation model and LiDAR depth measurements, respectively.
-At inference time, we wish to achieve rendered images for static $\{\hat{I}_{i}^{S},\hat{S}_{i}^{S}\}_{i=1}^{N}$ and dynamic $\{\hat{I}_{i}^{D},\hat{S}_{i}^{D}\}_{i=1}^{N}$ components of the scene for each timestamp independently, which can also be combined for a complete scene.
+RoDUS takes a set of $N$ input RGB images, their associated poses, and capturing timestamps $\{I_{i},p_{i},t_{i}\}_{i=1}^{N}$.
+Additionally, our model relies on semantic $\{S_{i}\}_{i=1}^{N}$, and geometric cues obtained from pre-trained 2D segmentation model and LiDAR depth measurements.
+At inference, we wish to achieve rendered images for static $\{\hat{I}_{i}^{S},\hat{S}_{i}^{S}\}_{i=1}^{N}$ and dynamic $\{\hat{I}_{i}^{D},\hat{S}_{i}^{D}\}_{i=1}^{N}$ components of the scene for each timestamp independently, which can also be combined for a complete view.
 As mentioned above, we study the quality of individual components rather than conforming solely to a satisfying composed result.
@@
 Preliminaries.
-RoDUS is based on NeRF [23] which is designed to capture both geometry and view-dependent appearance of a scene from a set of RGB images.
+RoDUS is based on NeRF [23], which is designed to capture both geometry and view-dependent appearance of a scene from a set of RGB images.
 It encodes the scene within a Multi-Layer Perceptron (MLP) and outputs color $\mathbf{c}(\mathbf{x},\mathbf{d})\in\mathbb{R}^{3}$ and density $\sigma(\mathbf{x})\in\mathbb{R}$ at any query position $\mathbf{x}\in\mathbb{R}^{3}$ along with its viewing direction $\mathbf{d}\in\mathbb{R}^{3}$.
@@
 Architecture.
-Inspired by [21, 18, 47], we factorize the scene into two elements, each learned through a separate pathway: one for moving components and the other for non-moving background.
+Inspired by [21, 18, 47], we factorize the scene into two elements, each learned through a separate pathway.
 Each branch is modeled independently using its own multi-level hash grid representation and MLPs.
-This approach enables the concurrent learning of both static and dynamic aspects of the scene.
-The static branch whose mission is to learn all non-moving parts in the scene, uses a 3D hash grid $\mathcal{H}^{S}$, analogous to the one used in [25].
+This approach enables the concurrent learning of both aspects of the scene.
+The static branch uses a 3D hash grid $\mathcal{H}^{S}$ for positions and positional embedding $\gamma(\cdot)$ for viewing directions, analogous to the one used in [25].
 To model lighting variations, we also condition on a per-frame latent embedding $l_{a}$ [21]:
 $(\sigma^{S},\mathbf{c}^{S},\mathbf{s}^{S})=\mathcal{F}^{S}_{\Theta}(\mathcal{H}^{S}(\mathbf{x}),\gamma(\mathbf{d}),l_{a}).$ (1)
-As the dynamic branch allows its outputs to depend on time, we employ a 4D hash grid $\mathcal{H}^{D}$ with one additional normalized timestamp $t$ as input.
-The dynamic field additionally incorporates a shadow head [47] designed to learn a shadow ratio $\rho\in[0,1]$.
+As the dynamic branch allows its outputs to depend on time, we extend to 4D hash grid $\mathcal{H}^{D}$ with one additional normalized timestamp $t\in[0,1]$ as input.
+We additionally incorporates a shadow head [47], designed to learn a shadow ratio $\rho\in[0,1]$.
 This scalar is later used to scale down the static radiance values $\mathbf{c}^{S}$ in Eq.
@@
 Figure 2: RoDUS’s architecture.
-Our model comprises two separate branches that take as input sampled positions $\mathbf{x}$, their viewing direction $\mathbf{d}$, and their timestamp $t$ for the dynamic branch, and generate outputs for every query coordinate.
+Our model comprises two separate branches that take as input sampled positions $\mathbf{x}$, viewing direction $\mathbf{d}$, and their timestamp $t$, and generate outputs for every query coordinate.
 Each branch is represented by a separate hash grid and the MLP-based neural function, which predicts colors, densities, and semantics.
-The rendered static image is used to calculate IRLS map $\omega(\epsilon)$ during the robust initialization step (Sec.
+The rendered static RGB is used to calculate IRLS map $\omega(\epsilon)$ during the robust initialization step (Sec.
 3.3), while dynamic semantic outputs are passed through a “foreground-only mask” to prevent over-explaining background regions (Sec.
 3.2).
-Rendering. We calculate the color $\hat{C}$ for each camera ray $\mathbf{r}$ by integrating all the sampled points along the ray:
+Rendering. We calculate the color $\hat{C}(\mathbf{r})$ of each ray $\mathbf{r}$ by integrating all the sampled points along the ray:
 $\hat{C}(\mathbf{r})=\sum_{i=1}^{K}T_{i}\left(\alpha_{i}^{S}(1-\rho_{i})\mathbf{c}_{i}^{S}+\alpha_{i}^{D}\mathbf{c}_{i}^{D}\right),$ (3)
 where $T_{i}=\exp\left(-\sum_{j=1}^{i-1}(\sigma_{j}^{S}+\sigma_{j}^{D})\delta_{j}\right)$ is the accumulated transmittance and $\alpha_{i}^{\xi}=1-\exp(-\sigma_{i}^{\xi}\delta_{i})$ with $\xi\in\{S,D\}$ are alpha values.
-Given the outputs of each branch, we can render static and dynamic maps separately.
+Given the outputs of each branch, we can also render static and dynamic maps separately.
 3.2 Enabling Semantic Awareness
@@
 It is shown in [34] that relying only on view-dependent radiance and sensitive mean square error makes the model prone to outliers and over-fit observations.
-On the other hand, semantic labels can be conceptualized as a view-invariant function [59], mapping world coordinates $\mathbf{x}$ to a distribution over semantic labels $\mathbf{s}(\mathbf{x})\in\mathbb{R}^{L}$.
-As a result, we choose to integrate a semantic head to predict the semantic class for each point, similar to the color head.
+On the other hand, semantic labels can be conceptualized as a view-invariant function [59], mapping world coordinates $\mathbf{x}$ to a distribution over semantic labels $\mathbf{s}(\mathbf{x})\in\mathbb{R}^{L}$, where $L$ is the number of classes.
+We choose to integrate a semantic head to predict the semantic class for each point, similar to the color head.
 Subsequently, the same rendering formula is applied for semantics to achieve 2D semantic segmentation maps:
@@
 In the forward pass (black arrows), the mask prevents the dynamic branch from outputting pixels that do not belong to foreground classes.
-Similarly, in the backward pass (orange arrows), it restricts the dynamic field from learning background pixels, which may primarily come from noisy annotations.
+While in the backward pass (orange arrows), it restricts the dynamic field from learning background pixels, which may primarily come from noisy annotations.
 (b) Predictions generated by 2D segmentation model are noisy and inconsistent between views.
@@
 Remarkably, this not only improves the stability of rendering semantics but also prevents half-precision floating point (float16) overflowing when working with frameworks such as tiny-cuda-nn [24].
-(a) Ground-truth
-(b) $\mathcal{L}_{robust}$
-(c) $\mathcal{L}_{robust}+\mathcal{L}_{sky}$
-(d) $\mathcal{L}_{robust}+\mathcal{L}_{road}$
-(e) $\mathcal{L}_{robust}+\mathcal{L}_{sky}+\mathcal{L}_{road}$
-Figure 4: (b) The use of a trimmed kernel helps completely remove moving cars in the scene.
-(c, d, e) As the sky and road regions are badly reconstructed, applying class constraints can aid the learning process of the static field.
+By performing inferences in 3D space and rendering to 2D with a foreground masking mechanism, our semantic outputs are multi-view consistent by design.
 3.3 Loss Formulation
@@
 Robust loss.
-Our objective is to reduce the sensitivity of the static model to dynamic objects by incorporating supervision through a robust photometric loss.
+Our objective is to reduce the sensitivity of the static model on dynamic components through a robust photometric loss.
 This loss can be formulated as an Iteratively Reweighted Least Squares (IRLS) problem:
@@
 hard weights).
-Furthermore, as every object has neighbor connections, they apply a spatial filter $\mathcal{F}$ on the residuals map to capture the smoothness, ensuring local support of the kernel:
+Furthermore, as every object has neighbor connections, we apply a spatial filter $\mathcal{F}$ on the residuals map to capture the smoothness, ensuring local support of the kernel:
 $\omega(\epsilon(\mathbf{r}))=\mathcal{F}\left[\epsilon(\mathbf{r})\leq\mathcal{T}_{\epsilon}\right],$ (6)
@@
 This filter aims to remove high-frequency details from being classified as outliers and allows them to be captured by the NeRF models.
+(a) Ground truth
+(b) $\mathcal{L}_{robust}$
+(c) $\mathcal{L}_{robust}+\mathcal{L}_{sky}$
+(d) $\mathcal{L}_{robust}+\mathcal{L}_{road}$
+(e) $\mathcal{L}_{robust}+\mathcal{L}_{sky}+\mathcal{L}_{road}$
+Figure 4: (b) Using a trimmed kernel effectively removes all moving cars from the scene.
+(c, d, e) As the sky and road regions are badly reconstructed, applying class constraints can aid the learning process of the static model.
 Sky loss.
@@
 $\begin{split}\mathcal{L}_{total}&=\underbrace{\mathcal{L}_{rgb}+\lambda_{d}\mathcal{L}_{d}+\lambda_{sem}\mathcal{L}_{sem}}_{reconstruction}\\ &+\underbrace{\mathcal{L}_{robust}+\lambda_{sky}\mathcal{L}_{sky}+\lambda_{road}\mathcal{L}_{road}}_{static\ regularization}+\underbrace{\lambda_{\sigma^{D}}\mathcal{L}_{\sigma^{D}}+\lambda_{\rho}\mathcal{L}_{\rho}+\lambda_{\mathcal{H}}\mathcal{L}_{\mathcal{H}}}_{dynamic\ regularization}\end{split}$ (9)
-3.4 Training strategy
-The use of robust estimation plays an important role in rejecting inconsistency (dynamic objects in our case) and guiding the static branch without any prior knowledge.
-However, we observe that using a binary kernel alone with a fixed threshold, after reaching a certain point throughout the training, the robust loss does not contribute to the overall loss anymore.
-We refer to this phenomenon as “kernel saturation”, this causes several parts of the scene not to be learned completely.
-Adjusting the threshold $\mathcal{T}_{\epsilon}$ seems to be a trivial solution, but given the fact that outdoor scenes are more complicated to handle, there is always a trade-off between the number of distractors removed and the static scene quality.
-Particularly, high-frequency but static details, equivalently appear as large residuals are considered as outliers, while low-frequency but moving regions such as shadows (dark) which mostly share a similar color with the road (gray) are uncontrollably included as inliers.
-Figure 4 portrays this, as there are still parts of the scene (e.g.
-the truck on the right) that are not fully learned by the model.
-In contrast, relying solely on reconstruction losses leads to several local optima during optimization due to the entanglement of geometry and appearance.
-We propose using the robust loss as an initialization step for the static model to learn the basic overall structures of the scene before delving into complete scene understanding.
-The initialization step guides the network to prioritize consistent (or confident) regions that further help stabilize the training process.
-Therefore, we start the model training with the robust loss (Eq.
-5) and sky loss (Eq.
-7) enabled for the first epoch, and subsequently enable the remaining losses.
-This serves as a foundation for disentangling the scene in later training stages.
 GT
@@
 Figure 5: We demonstrate RoDUS’s capabilities on multiple tasks.
-Notice that we successfully retain separation consistencies between all photometric, geometric, and semantic outputs.
 Legend: GT: Ground truth, $\hat{I}$: Composed RGB, $\hat{d}$: Composed depth, $I^{S}$: Static RGB, $\hat{d}^{S}$: Static depth, $\hat{I}^{D}$: Dynamic RGB (blended with a green background), $\hat{S}$: Composed segmentation map, $\hat{S}^{S}$: Static segmentation map, $M^{D}$: Motion segmentation.
 $M^{D}$
+3.4 Training Strategy
+The use of robust estimation plays an important role in rejecting inconsistency (dynamic objects in our case) and guiding the static branch without any prior knowledge.
+However, we observe that when using a binary kernel alone with a fixed threshold, after reaching a certain point throughout the training, the robust loss does not contribute to the overall loss anymore.
+We refer to this phenomenon as “kernel saturation”, this causes several parts of the scene not to be learned completely.
+Adjusting the threshold $\mathcal{T}_{\epsilon}$ seems to be a trivial solution, but given the fact that outdoor scenes are more complicated to handle, there is always a trade-off between the number of distractors removed and the static scene quality.
+Particularly, high-frequency but static details, equivalently appear as large residuals are considered as outliers, while low-frequency but moving regions such as shadows (dark) which mostly share a similar color with the road (gray) are uncontrollably included as inliers.
+Fig.
+6 portrays this, as there are still parts of the scene (e.g.
+road marking) that are not fully learned.
+In contrast, relying solely on reconstruction losses leads to several local optima during optimization due to the entanglement of geometry and appearance.
+We propose using the robust loss as an initialization step for the static model to learn the basic overall structures of the scene before delving into complete scene understanding.
+The initialization step guides the network to prioritize consistent (or confident) regions that further help stabilize the training process.
+Therefore, we start the model training with the robust loss (Eq.
+5) and sky loss (Eq.
+7) enabled for the first epoch, and subsequently enable the remaining losses.
+By doing so, we can confidently select a low threshold without compromising the quality.
+This serves as a foundation for disentangling the scene in later training stages.
 4 Evaluation
@@
 We refer to the supplementary for additional details on the selection of the training sequences and data pre-processing.
-[…]
+Baseline.
+We benchmark RoDUS against 4 different state-of-the-art methods for static scene extraction (RobustNeRF [34]) and dynamic scene rendering (D2NeRF [47], SUDS [43], and EmerNeRF [52]).
+Evaluations focus on two key factors: the quality of novel view synthesis of the static image and the segmentation masks of dynamic objects.
+While our primary focus is on static-dynamic disentanglement, RoDUS’s design allows it to represent dynamic scenes holistically so the quality of the overall image is also considered.
+We showcase the performance of our method on three representative scenarios in Fig.
+5
 Table 1: Quantitative comparisons with SoTA methods on KITTI-360 and Pandaset.
@@
 […]
-KITTI-360 RobustNeRF [${}^{{\dagger}}$ 19.50 0.722 0.282 - - -
-[…]
+KITTI-360 RobustNeRF [34] - - - 19.50 0.722 0.282 - - -
+D2NeRF [47] 23.86 0.688 0.232 18.70 0.588 0.309 12.04 4.67 7.41
 SUDS [43] 28.16 0.836 0.175 18.81 0.687 0.298 93.89 5.35 9.34
@@
 RoDUS (ours) 26.47 0.857 0.148 23.43 0.802 0.181 79.45 64.84 76.52
-Pandaset RobustNeRF [${}^{{\dagger}}$ 23.41 0.764 0.197 - - -
-[…]
+Pandaset RobustNeRF [34] - - - 23.41 0.764 0.197 - - -
+D2NeRF [47] 25.25 0.692 0.291 23.38 0.682 0.292 - - -
 SUDS [43] 30.28 0.851 0.138 25.34 0.788 0.182 - - -
@@
 RoDUS (ours) 28.39 0.877 0.135 25.83 0.825 0.156 - - -
-${}^{{\dagger}}$ As RobustNeRF rejects distractors during training, full image reconstruction metrics are not included.
-GT
-RobustNeRF [34]
-N/AN/A
-[…]
-SUDS [43]
-EmerNeRF [52]
-RoDUS (ours)
-Composed RGB
-Static RGB
-Dynamic RGB
-[…]
+As RobustNeRF rejects distractors during training, full image reconstruction metrics are not included.
+GT Refer to caption
+RobustNeRF [34] N/A Refer to caption N/A
+D2NeRF [47] Refer to caption Refer to caption Refer to caption
+SUDS [43] Refer to caption Refer to caption Refer to caption
+EmerNeRF [52] Refer to caption Refer to caption Refer to caption
+RoDUS (ours) Refer to caption Refer to caption Refer to caption
+Composed RGB Static RGB Dynamic RGB
+Figure 6: Reconstruction and decomposition qualitative results comparisons on Pandaset.
+D2NeRF struggles in intricate driving contexts and produces poor results.
+SUDS faces difficulties in distinguishing between ego-motion and object-motion due to its reliance on optical flow supervision, leading to a large number of cloudy artifacts in the dynamic image to match the ground truth optical flows.
+RobustNeRF successfully reconstructs the background with no dynamic objects remaining, but it suffers from saturation, leading to degraded results (as explained in Sec.
+3.3).
+We can see that the separation quality is still suboptimal.
+In contrast, RoDUS successfully decouples dynamic objects and achieves a clean background with no artifacts remaining.
 4.2 Static-Dynamic Decomposition
@@
 4.3 Dynamic Objects Segmentation
-[…]
-GT image Refer to caption Refer to caption Refer to caption
+For this task, we obtain the motion mask $M^{D}$ by rendering dynamic opacities and applying a threshold $\mathcal{T}_{O^{D}}=0.5$.
+We report mask Recall, IoU, and F1-score on training views.
+We only evaluate on KITTI-360 due to the absence of ground-truth instance masks in Pandaset.
+In Fig.
+7, we explore various scenarios to understand how the ego-actor moves in relation to other moving objects and how it observes the scenes that could impact the decomposition.
+Particularly, case (a) is likely the easiest as the background is broadly observed for a long period, we can still notice a good separation in D2NeRF, SUDS’s results.
+In cases (b) and (c), when the ego actor is in motion, most methods suffer from incorrect segmentation, while our semantic guidance forces the model to learn the correct dynamic objects.
+SUDS tends to produce large dynamic areas, resulting in very high Recall but low IoU and F1, yet EmerNeRF exhibits sensitivity to lighting variations.
+Correspondingly, Tab.
+1 indicates that RoDUS’s metrics outperform the others in both tasks.
+GT Refer to caption Refer to caption Refer to caption
 GT mask Refer to caption Refer to caption Refer to caption
-[…]
+D2NeRF [47] Refer to caption Refer to caption Refer to caption
 SUDS [43] Refer to caption Refer to caption Refer to caption
@@
 8 for visual comparisons.
+Table 2: Ablation study. Impact of our design choice on static NVS task.
 […]
-Refer to caption w/o $\mathcal{L}_{road}$ Refer to caption w/o $\mathcal{L}_{robust}$ Refer to caption Full model static RGB Refer to caption Full model static depth Figure 8: Static depth map ablation
+Full model 25.94 0.814 0.170
+w/o $\mathcal{L}_{d}$0.80 -0.032 +0.023
+w/o $\mathcal{L}_{sem}$1.67 -0.080 +0.078
+w/o $\mathcal{L}_{robust}$0.91 -0.047 +0.067
+w/o $\mathcal{L}_{road}$0.84 -0.035 +0.026
+w/o $\mathcal{L}_{\sigma^{D}}$0.74 -0.030 +0.024
+w/o FG mask -0.67 -0.035 +0.023
+w/o FG mask
+w/o $\mathcal{L}_{sem}$
+w/o $\mathcal{L}_{robust}$
+Full model
+Figure 8: Components ablation on $\hat{I}^{S}$.
 4.5 Limitation
-Our approach can’t handle areas that lack observations or have been highly occluded most of the time within the sequence.
-This is particularly notable by large or slow-moving objects, illustrated in Fig.
+In general, RoDUS works best with scenes where labels belonging to potentially movable objects are well-defined.
+Nonetheless, if the semantic ground truth is absent or heavily mislabeled, it moderately impacts performance.
+Moreover, our approach can’t handle areas that lack observations or have been highly occluded most of the time within the sequence.
+This is particularly notable by large or slow-moving objects in Fig.
 9.
-Exploring a model that is capable of detecting and inpainting these missing regions based on prior knowledge will be interesting as a future research direction.
+Exploring a model that is capable of uncertainty modeling to incorrect or anomalous pixels and inpainting these missing regions based on prior knowledge will be interesting as a future research direction.
 GT
@@
 In this paper, we presented RoDUS, a method for disentangling dynamic scenes using neural representation.
-Our approach effectively separates moving components from the scene without the need for manual labeling.
+Our approach effectively separated moving components from the scene without the need for manual labeling.
 To enhance disentangling efficiency, we proposed a robust training strategy and incorporated semantic reasoning, resulting in high-quality background view synthesis and accurate segmentation.
-We conducted extensive experiments on challenging autonomous driving datasets, showcasing that RoDUS surpasses current state-of-the-art NeRF frameworks.
+Extensive experiments on challenging autonomous driving datasets showcased that RoDUS surpasses current state-of-the-art NeRF frameworks.
 Appendix 0.A Implementation Details
 0.A.1 Data Processing
-In our experiments, we downsample Pandaset images by a factor of 4 ($480\times 270$) and KITTI-360 by a factor of 2 ($704\times 188$).
+In our experiments, we downsample Pandaset images by a factor of 4 ($480\times 270$) and KITTI-360 images by a factor of 2 ($704\times 188$).
 The selected scenes from KITTI-360 are summarized in Tab.
 3.
-For Pandaset, we used sequences 023, 042, and 043.
-We specifically selected those scenes because they feature multiple dynamic objects while providing sufficient views of the background regions.
+For Pandaset, we use sequences 023, 042, and 043.
+We specifically select those scenes because they feature multiple dynamic objects while providing sufficient views of the background regions.
 Table 3: Selected KITTI-360 sequences for training, with skip step of 2.
@@
 Our hash grids have the size of $2^{19}$ for $16$ levels with the base resolution and finest resolution of 16 and 2048 respectively.
-Additionally, our dynamic hash grid allows 4D inputs, which means that the intermediate feature vector is computed through a quadrilinear interpolation of 16 grid point vectors instead of the 8 used for the 3D hash grid.
+Additionally, our dynamic hash grid allows 4D inputs, which means that the intermediate feature vector is computed through a quadrilinear interpolation of 16 grid point vectors instead of 8 used in the 3D hash grid.
 The static branch of our model is equipped with a 16-dimension per-frame embedding to adapt to varying lighting conditions.
@@
 0.A.4 Training
-We trained RoDUS for 50 epochs using the Adam optimizer with a batch size of 200.
-The learning rate was set to $0.01$, and we used cosine decay to decrease it to $0.0001$.
+We train RoDUS for 50 epochs using the Adam optimizer with a batch size of 200.
+The learning rate is set to $0.01$, and decayed to $0.0001$.
 The rays are grouped in $15\times 15$ patches to suit DSSIM, planar regularization, and robust losses.
-Section 0.A.4 summarizes the hyperparameters used for the training of RoDUS, we do not require manual tuning for each scene, but we believe that such tuning could potentially further enhance performance.
-Table 4: Training hyperparameters
+Sec.
+0.A.4 summarizes the hyperparameters used for the training of RoDUS, we do not require manual tuning for each scene, but we believe that such tuning could potentially further enhance performance.
+Table 4: Training hyperparameters.
 Parameter Value
@@
 $\lambda_{\mathcal{H}}$ 0.01
-[…]
+Baseline Implementations.
+We adapt the code in the official repository of each method: RobustNeRF [^1] https://robustnerf.github.io/, D2NeRF [^2] https://github.com/ChikaYan/d2nerf, SUDS [^3] https://github.com/hturki/suds, EmerNeRF [^4] https://github.com/NVlabs/EmerNeRF and modify their dataloaders.
+We search for the best robust loss threshold $\mathcal{T}_{\epsilon}$ for RobustNeRF, while the rest are trained with the default parameters provided in the repositories.
 Appendix 0.B Additional Results
@@
 Furthermore, RoDUS can be trained from noisy semantic annotations, leading to a scene-consistent semantic representation.
-A combination of multiple modalities also supports the model in reconstructing and segmenting thin structures and objects (e.g.
+Our combination of multiple modalities also supports the model in reconstructing and segmenting thin structures and objects (e.g.
 light poles, and traffic signs).
-GT
-RobustNeRF [34]
-N/AN/A
-[…]
-SUDS [43]
-EmerNeRF [52]
-Figure 10: Additional qualitative comparisons on Pandaset.
-RoDUS (ours)
-Composed RGB
-Static RGB
-Dynamic RGB
 GT Refer to caption
-RobustNeRF [34]
-N/A
-N/A
-[…]
+RobustNeRF [34] N/A Refer to caption N/A
+D2NeRF [47] Refer to caption Refer to caption Refer to caption
 SUDS [43] Refer to caption Refer to caption Refer to caption
@@
 RoDUS (ours) Refer to caption Refer to caption Refer to caption
-GT image Refer to caption
-RobustNeRF [34]
-N/A
-N/A
-[…]
+Composed RGB Static RGB Dynamic RGB
+Figure 10: Additional qualitative comparisons on Pandaset.
+GT Refer to caption
+RobustNeRF [34] N/A Refer to caption N/A
+D2NeRF [47] Refer to caption Refer to caption Refer to caption
+SUDS [43] Refer to caption Refer to caption Refer to caption
+EmerNeRF [52] Refer to caption Refer to caption Refer to caption
+RoDUS (ours) Refer to caption Refer to caption Refer to caption
+GT Refer to caption
+RobustNeRF [34] N/A Refer to caption N/A
+D2NeRF [47] Refer to caption Refer to caption Refer to caption
 SUDS [43] Refer to caption Refer to caption Refer to caption
@@
 $\hat{S}^{S}$
-Figure 12: Additional results on RoDUS’s capability.
+Figure 12: Additional showcases of RoDUS’s capability.
 $M^{D}$
@@
 With $\mathcal{T}_{\epsilon}=0.7$, there are still a lot of structures that could not be learned (e.g.
-the far away building), when we increase $\mathcal{T}_{\epsilon}$ to $0.9$, the shadow regions started to appear.
-As a result, we choose $\mathcal{T}_{\epsilon}=0.75$ as it yields the best visual quality with the least artifacts.
-Correspondingly, the kernel saturates as we observed no improvement as the training continues.
-Notice that the model could not learn the road marking in any case.
+the far away building), when we increase $\mathcal{T}_{\epsilon}$ to $0.9$, the shadow regions start to appear.
+As a result, we select $\mathcal{T}_{\epsilon}=0.75$ for benchmarking as it yields the best visual quality with the least artifacts.
+Correspondingly, the kernel saturates and we observed no improvement as the training continues.
+Notice that the model still could not learn the road marking in any case.
 GT
@@
 Figure 13: Qualitative comparison on static RGB image with different threshold values.
-0.C.2 Shadow Handling
+0.C.2 Various Challenging Scenarios
+Fig.
+14 illustrates RoDUS’s performance in low light (driving at night) and uneven illumination (driving into a shaded area) environments with high variations between views which are common in an autonomous driving context.
+It also highlights that the model could handle chaotic traffic and slow-moving entities (i.e.
+cyclists) as long as there is no significant occlusion.
+Complex traffic Refer to caption Refer to caption Refer to caption
+Cyclist Refer to caption Refer to caption Refer to caption
+Low light Refer to caption Refer to caption Refer to caption
+Uneven Illumination Refer to caption Refer to caption Refer to caption
+Composed RGB Static RGB Dynamic RGB
+Figure 14: RoDUS’s adaptivity on different scenarios.
+0.C.3 Shadow Handling
 Shadows cast by dynamic objects are also time-varying and modeled separately in our model. Fig. 15 shows that RoDUS successfully decouples the shadows from the scene.
@@
 We believe this area of study could hold significant potential for advancements in the field, providing opportunities to enhance the robustness and security of digital content.
-0.D.2 Limitation
+0.D.2 Limitations
 Apart from the main limitation mentioned in the paper, we also encountered several problems that are related to most NeRF-based methods in general: RoDUS demands precise camera poses, which require multiple sources of odometry when working with outdoor, high-dynamic scenes.